The length of a rectangle is eight-fifths times the width. The are is 160. What are the dimensions of the rectangle?
Answer by unlockmath(1688)

Hello,
We can set up the following: Let x be the width and 8/5x be the length so the equation will look like:
x*8/5(x)=160
Rewritten as:
8/5x^2=160
Multiply by 5 to get:
8x^2=800
Divide by 8 and square root both sides to get:
x=10 is the width
16 is the length.
Make sense?
RJ
www.math-unlock.com